Carla Company, which is subject to a 40% income tax rate, projected its income before taxes for next year as shown here:
Sales (184,000 units) $9,200,000
Cost of sales
Variable costs 2,300,000
Fixed costs

3,450,000

Pretax earning

$3,450,000

If Carla wants $5,175,000 in pretax earning, what is the required level of sales, in dollars?
Total Revenue
Answer #1

Contribution margin=Sales-Variable cost

=(9,200,000-2,300,000)=$6,900,000

Contribution margin ratio=Contribution margin/Sales

=6,900,000/9,200,000

=0.75

Target Contribution margin=Fixed cost+Target profits

=(3,450,000+5,175,000)=$8625000

Hence required sales=8625000/0.75

=$11,500,000.